OpenStudy (anonymous):

A farmer wishes to enclose a pasture that is bordered on one side by a river (so one of the four sides wont require fencing) She has decided to create a rectangular shape for the area and will use barbed wire to create the enclosure there are 600 feet of wire availible for this project and she will use all the wire. what is the maximum area enclosed by the fence then find the maximum of the function.

OpenStudy (anonymous):

The area enclosed is (2x)(600-2x). \[(2x)(600-2x)=1200x-4x^{2}\] In order to maximize we need the derivative to be equal to 0. \[y'=1200-8x=0\] \[1200 = 8x\] \[x=150\] Therefore the sides for maximum area are 150*300.

OpenStudy (anonymous):

The area is: 45000

OpenStudy (anonymous):

whats the maximum of the functions?

OpenStudy (radar):

Let the sides that are perpendicular to the river be x feet in length Then the remaining side will be 600-2x in length. (Perimeter = 2x +(600-2x) Area is length times width or x(600-2x)=600x-2x^2 A=600x-2x^2 dy/dx=-4x+600 set this to = 0 for maxima/minima -4x+600=0 -4x=-600 x=150 so the dimensions for maximum area or x=150 and 600-(2 * 150) =300 all in feet of course.

OpenStudy (radar):

You solved for the maximum. x=150

OpenStudy (phi):

Hi kassia, you are doing pre-calc, right? So "they" want you to solve this by setting up an equation that represents a parabola, and you say, "OK, the parabola peaks at its vertex, let me find the vertex" Easier said then done...

OpenStudy (anonymous):

yes phi you read my mind

OpenStudy (anonymous):

If you're doing pre-calc I apologize. It read like a Calc-1 question I've had before :)

OpenStudy (phi):

Here's how you start. You have four sides, it's a rectangle, so you opposite sides are equal. So the total perimeter is, for example, x+y+x+y, where x and y are, say, the length and width.

OpenStudy (phi):

The equation for the perimeter is 2x + 2y = ?? Help me here, please.

OpenStudy (phi):

Meanwhile, here's an equation for the area of the rectangle. Area= xy Area is length times width.

OpenStudy (radar):

The amount of fencing is to be used only on three sides and the amount of fencing to be used is 600 feet. this information is to be used to solve this problem.

OpenStudy (phi):

Oops, thanks. Maybe kassia will fix my mistake??

OpenStudy (anonymous):

hey sorry the computer im on is actually retarded but wouldnt it be 2x+y=?? because one of the sides water

OpenStudy (anonymous):

hey sorry the computer im on is actually retarded but wouldnt it be 2x+y=?? because one of the sides water

OpenStudy (radar):

I have to run, but I did work it out step by step for kassia.

OpenStudy (anonymous):

hey sorry the computer im on is actually retarded but wouldnt it be 2x+y=?? because one of the sides water

OpenStudy (anonymous):

kjkl

OpenStudy (anonymous):

kjkl

OpenStudy (anonymous):

kjkl

OpenStudy (anonymous):

kjkl

OpenStudy (phi):

Yes, 2x + y= ?, but the next step is what goes in the place of the question mark? we want an equation. hint, we are very close.

OpenStudy (phi):

The delay is annoying. So let's forge ahead.... the equation for the perimeter is 2x+y= 600 Area= x y Next step is to replace y in the Area equation, so there are only x's. Why, well because you have done so many of these problems you just know!

OpenStudy (phi):

Hi Kas, Did you give up? Or did you just want the answer w/o knowing all the details?

OpenStudy (anonymous):

oh sorry i got distracted alright so should i plug in 3 for y?

OpenStudy (phi):

Solve for y using the perimeter equation: 2x+y= 600 y= ??

OpenStudy (anonymous):

but i cant solve for y if i dont have x

OpenStudy (phi):

By solve, I mean just use algebra so y is by itself on one side of the equal sign, and everything else is on the other side. True we don't know what x is,but that's ok.

OpenStudy (anonymous):

so y=300-x or y=600-2x

OpenStudy (phi):

To solve for y, subract 2x from both sides 2x+y-2x= 600 - 2x On the left, the 2x-2x cancel, so you get y= 600 - 2x So your 2nd answer is good, the first not so good. As a boring aside, notice if we start with y= 600-2x and divide both sides by 2 you get y/2 = 300 -x (but that's not what we want. Next step, in the equation Area= x y replace y with its equivalent 600-2x. Then multiply it out.

OpenStudy (anonymous):

\[2x^2+600x\]

OpenStudy (phi):

So you should get a new equation Area= x(600-2x) Area = 600x - 2 x^2 or, rearranging \[A= -2x ^{2}+600x\] Did you get this?

OpenStudy (anonymous):

yeah so then should i take out a 2x so i get 2x(-x+300) and then can i just set both equations equal to zero?

OpenStudy (phi):

factor out 2x would be going backwards. What you should be doing is looking at the equation for area A= -2 x^2 + 600x and it should remind you of the equation for a parabola (because surprise!) it is.

OpenStudy (anonymous):

okay so whats the next step then?

OpenStudy (phi):

First, I just noticed you left off the minus sign off x^2 I hope that was just a typo. other wise you'll have to practice multiplying. The trouble with math is a tiny mistake ruins everything. OK, you have an equation that tells you the area if you know one side of the fence (labelled x). So the area you get if the side has zero length is 0. if x=1, Area is -2*1^2 +600(1) = -2 + 600 = 598 You get the idea? But we know this is a parabola, and it has a peak at its vertex A , the area, peaks at the parabola's vertex.

OpenStudy (phi):

So the next step is to find the x value of the vertex. Do you remember from yesterday? the vertex is -b/(2a) for the parabola y=a x^2 +b x +c

OpenStudy (phi):

Did I confuse you??

OpenStudy (phi):

Once you find the x value of the vertex, use that x value in the equation for the Area Area= -2 x^2 + 600 x That's going to be the maximum area you can enclose with your fence.

OpenStudy (phi):

If everything works out, we should get the same answer as from Calculus.

OpenStudy (anonymous):

got it thanks!

OpenStudy (phi):

Excellent! Can you post your work?